T heta1.eps cuarto.eps

T heta2.eps cuarto.eps

T heta3.eps cuarto.eps

T heta4.eps cuarto.eps

plugin_{T} heta1.epsEstimacinplug − indeladensidaddelprimerperiodode2010 plugin_{T} heta2.epsEstimacinplug − indeladensidaddelsegundoperiodode2010 plugin_{T} heta3.epsEstimacinplug − indeladensidaddeltercerperiodode2010 plugin_{T} heta4.epsEstimacinplug − indeladensidaddelcuartoperiodode2010 T heta1.epsV alidacincruzadadelprimerperiodode2010 T heta1.epsEstimacinL2deladensidaddelprimerperiodode2010 T heta1.epsEstimacinKLdeladensidaddelprimerperiodode2010 v_{T} heta1.epsEstimacinnoparamtricadelaDensidadKLdelprimerperiodode2010.v = 50(azul), v = 100(verde), v = 150(violeta) no_{p}aramT heta1.epsEstimacinnoparamtricadelaDensidad.L2(azul), KL(verde), plug − in(violeta)

T heta2.epsV alidacincruzadadelsegundoperiodode2010 T heta2.epsEstimacinL2deladensidaddelsegundoperiodode2010 T heta2.epsEstimacinKLdeladensidaddelsegundoperiodode2010 v_{T} heta2.epsEstimacinnoparamtricadelaDensidadKLdelsegundoperiodode2010.v = 50(azul), v = 100(verde), v = 150(violeta) no_{p}aramT heta2.epsEstimacinnoparamtricadelaDensidad.L2(azul), KL(verde), plug − in(violeta) T heta3.epsV alidacincruzadadeltercerperiodode2010 T heta3.epsEstimacinL2deladensidaddeltercerperiodode2010 T heta3.epsEstimacinKLdeladensidaddeltercerperiodode2010 v_{T} heta3.epsEstimacinnoparamtricadelaDensidadKLdeltercerperiodode2010.v = 50(azul), v = 100(verde), v = 150(violeta) no_{p}aramT heta3.epsEstimacinnoparamtricadelaDensidad.L2(azul), KL(verde), plug − in(violeta)

T heta4a.epsV alidacincruzadadelcuartoperiodode2010 T heta4.epsEstimacinL2deladensidaddelcuartoperiodode2010 T heta4.epsEstimacinKLdeladensidaddelcuartoperiodode2010 v_{T} heta4.epsEstimacinnoparamtricadelaDensidadKLdelcuartoperiodode2010.v = 50(azul), v = 100(verde), v = 150(violeta) no_{p}aramT heta4.epsEstimacinnoparamtricadelaDensidad.L2(azul), KL(verde), plug − in(violeta) κ = 1/h_{s}^{2} regresionhsB1.epsEstimacindelafuncinderegresincircularlinealenB1.Estimacindelparmetrodesuavizadomediante_{s}^{2} regresionhsB2.epsEstimacindelafuncinderegresincircularlinealenB2.Estimacindelparmetrodesuavizadomediante_{s}^{2} regresionhsC9.epsEstimacindelafuncinderegresincircularlinealenC9.Estimacindelparmetrodesuavizadomediante_{s}^{2} regresionhsF 2.epsEstimacindelafuncinderegresincircularlinealenF 2.Estimacindelparmetrodesuavizadomediante_{s}^{2} regresionhsG2.epsEstimacindelafuncinderegresincircularlinealenG2.Estimacindelparmetrodesuavizadomediante_{s}^{2} corrT heta1.epsCoeficientedecorrelacinparaladireccindelvientodel2010(1/4) plotT heta1.epsGrficadelaseriedetiempodel2010(1/4)frentesusretardos ts_{k}T heta1.epsSeriedetiempodel2010(1/4) corrT heta2.epsCoeficientedecorrelacinparaladireccindelvientodel2010(2/4) ts_{k}T heta2.epsSeriedetiempodel2010(2/4) corrT heta3.epsCoeficientedecorrelacinparaladireccindelvientodel2010(3/4) plotT heta3.epsGrficadelaseriedetiempodel2010(3/4)frentesusretardos ts_{k}T heta3.epsSeriedetiempodel2010(3/4) corrT heta4.epsCoeficientedecorrelacinparaladireccindelvientodel2010(4/4)